### v1

@@ Abstract
We investigate the stability of continuous-time quantum walks (CTQWs) in a range of network topologies under different decoherence mechanisms, defining stability as the system’s ability to preserve quantum properties over time. The networks studied range from homogeneous to heterogeneous structures, including cycle, complete, Erdős–Rényi, small-world, scale-free, and star topologies. The decoherence models considered are intrinsic decoherence, Haken–Strobl noise, and quantum stochastic walks (QSWs). To assess quantum stability, we employ several metrics: node occupation probabilities, the $\ell_{1}$-norm of coherence, fidelity with the initial state, quantum-classical distance, and von Neumann entropy. Our results reveal that the interplay of both network topology and decoherence model influences coherence preservation. Intrinsic decoherence results in the slowest decay of coherence, followed by Haken–Strobl noise, while QSW causes the most rapid loss of coherence. The stability ranking among network topologies varies depending on the decoherence model and quantifier used. For example, under Haken–Strobl and intrinsic decoherence, the quantum-classical distance ranks the cycle network more stable than scale-free networks, although other metrics consistently favour scale-free topologies. In general, heterogeneous networks, such as star and scale-free networks, exhibit the highest stability, whereas homogeneous topologies, such as cycle and Erdős–Rényi networks, are more vulnerable to decoherence. The complete graph, despite its homogeneity, remains highly stable due to its dense connectivity. Furthermore, in heterogeneous networks, the centrality of the initialized node, measured by degree or closeness, has a pronounced impact on stability, underscoring the role of local topological features in quantum dynamics.
keywords:
@@
Continuous-time Quantum Walks, Complex Networks, Decoherence Dynamics, Quantum Coherence, Quantum Stability
1 Introduction
@@
Quantum walks (QWs), the quantum analogs of classical random walks, have emerged as a powerful framework for exploring a wide range of quantum phenomena, including quantum computation, coherent transport, and the dynamical behavior of complex quantum systems [1, 2, 3]. Unlike their classical counterparts, QWs exploit quantum superposition and interference, enabling the walker to explore graphs more efficiently and, in some cases, achieve quantum speedups in algorithmic tasks [4].
@@
QWs can be broadly categorized into two types: discrete-time quantum walks (DTQWs) [1] and continuous-time quantum walks (CTQWs) [5]. DTQWs evolve via alternating applications of a coin operator and a conditional shift, allowing fine-grained control over the walker’s trajectory [6]. This approach has proven useful in quantum search algorithms [7] and in simulating topologically nontrivial phases [8]. In contrast, CTQWs evolve under a time-independent Hamiltonian defined by the graph structure, without requiring an internal coin degree of freedom. This formulation is particularly well-suited for modeling coherent transport in structured networks and has applications in quantum search [4, 9, 10], quantum state transfer [11], entanglement routing [12], and energy transport in photosynthetic complexes [13].
@@
However, real-world implementations of QWs are inevitably affected by decoherence arising from environmental interactions and imperfections in the system. These effects disrupt unitary evolution, suppress quantum coherence, and reduce transport efficiency and computational fidelity [14, 15]. Interestingly, in some cases decoherence can also assist transport [16, 17]. Understanding how decoherence influences QWs is therefore critical for the development of robust quantum technologies.
@@
Since decoherence in open quantum systems is typically continuous rather than discrete [14], our study focuses on CTQWs under a variety of continuous decoherence mechanisms. Common modeling approaches include intrinsic decoherence, modeled as dephasing in the energy eigenbasis via the Milburn equation [18]; position-basis decoherence, described by the Haken–Strobl master equation [19]; and quantum stochastic walks (QSW), which interpolate between quantum and classical regimes [20]. These frameworks thus enable us to explore different facets of open quantum dynamics and their consequences for quantum transport.
@@
The evolution of CTQWs is influenced by both the form of decoherence and the structure of the underlying graph [21, 22]. Earlier studies have explored CTQW dynamics on one-dimensional lattices to examine quantum transport [23] and quantum correlations [24], while more recent efforts have extended this work to symmetric and complex networks, assessing how structure interacts with decoherence [21, 25, 26].
@@
In this work, we broaden the scope by considering a diverse set of network topologies, from homogeneous to highly heterogeneous, and evaluating their behavior under multiple decoherence models. Our primary objective is to understand how the interplay between network structure and environmental noise governs the behavior and stability of CTQWs.
@@
Motivated by analogous findings in classical network theory, where the stability of dynamical states is often determined by the interplay between topology and dynamics [27, 25, 28], so in this study we seek to uncover similar principles within the context of quantum systems. Specifically, we investigate how the CTQW dynamics evolves under various decoherence mechanisms across different network architectures. Here, we define quantum stability as the ability of a system to retain its quantum features, such as coherence, over time despite the presence of decoherence.
@@
It is well established that node’s centralities, such as degree, betweenness, and closeness, play a critical role in determining the resilience of classical networks [29, 30]. Recent studies indicate that these structural attributes also significantly influence quantum dynamics [31].Identifying which topological features help mitigate or exacerbate decoherence is essential for designing stable and efficient quantum systems. Such insights are crucial for advancing robust quantum communication protocols, quantum devices, and fault-tolerant quantum architectures [32]. Motivated by this, our study also investigates how node centralities affect quantum stability, with a particular focus on heterogeneous networks, where centrality disparities are pronounced. To quantify the stability of CTQWs under different decoherence models, we employ several metrics commonly used in the analysis of open quantum systems, including quantum-classical distance, von Neumann entropy, $\ell_{1}$-norm of coherence, and fidelity. These measures provide a multifaceted view of how far the system deviates from ideal quantum behavior over time. For example, the quantum-classical distance has been used to study how CTQWs diverge from quantum behavior under decoherence in simple network topologies such as complete, cycle, and star graphs [21].
@@
We begin Sec. 2 by detailing the dynamics of CTQWs in the presence of various decoherence mechanisms. In Sec. 3, we provide an overview of the network structures considered in our study. Sec. 4 introduces the quantitative measures employed to assess the stability of CTQWs across different graph topologies. The primary findings of our investigation are presented in Sec. 5, where we analyze how decoherence impacts quantum transport on both simple and complex networks. Finally, Sec. 6 concludes the paper with a summary of our key results and their broader implications.
2 Dynamical framework of quantum networks
@@
The Laplacian matrix is defined as $L=D-A$, where $D$ is the diagonal degree matrix with the degree (i.e., number of connections) of each node as its diagonal elements, and $A$ is the adjacency matrix, with elements $A_{ij}=1$ if node $i$ and $j$ are connected and $A_{ij}=0$ otherwise.
We investigate the impact of three distinct types of decoherence on CTQW: (i) intrinsic decoherence, modeled by the Milburn equation [18], (ii) position-basis decoherence, described by the Haken–Strobl master equation [19], and (iii) quantum stochastic walks, as formulated in [20]. Each of these models introduces a unique form of non-unitary evolution, enabling us to explore different facets of quantum system dynamics under decoherence on networks.
In the absence of decoherence, that is, ideal noiseless conditions, the quantum walker evolves coherently, governed by unitary dynamics. This evolution is described by the Liouville–von Neumann equation [33]:
@@
To model intrinsic decoherence, we take Milburn formalism [18], which introduces non-unitary dynamics by modifying the noiseless evolution Eq. (1) as follows:
@@
Here, $\gamma$ represents the decoherence rate. This model captures intrinsic decoherence, which reflects the system’s internal loss of coherence, independent of external environmental interactions. In our simulations, we set $\gamma=0.1$ without loss of generality, following [21]. This value introduces reasonable decoherence effects while preserving essential features of quantum dynamics. The double-commuter term in Eq. (2) models loss of coherence in the energy basis.
The next model we consider is based on the Haken-Strobl master equation, which is commonly employed to describe dephasing noise on the position basis. It takes the following form [19]:
@@
$P_{k}=|k\rangle\langle k|$ are projectors on the position basis. This equation models decoherence by coupling the system to a bath that causes loss of coherence between different positions without energy exchange.
The last decoherence model is Quantum Stochastic Walk (QSW) [20] that interpolates between pure quantum and classical behavior. It is described by the Gorini–Kossakowski–Sudarshan–Lindblad master equation [34, 35, 36] as follows:
@@
$P_{kj}=L_{kj}|k\rangle\langle j|$ and $p$ is the parameter that controls the interpolation between unitary evolution and stochastic noise. We consider $p=0.1$, which corresponds to small decoherence, but previous studies [21] have shown that it is enough to study the effect of quantum stochastic decoherence.
3 Network Topologies
We consider a quantum walker evolving under CTQW dynamics, where the walker is initially localized at a single node in the network. This approach is applied consistently across all the networks that we consider in this study. The rationale behind this choice is to minimize the fidelity between the quantum-evolved state and the classically evolved state, as discussed in prior studies on CTQW [37]. As the system evolves, the walker’s state becomes a quantum superposition over multiple nodes, with the probability of finding the walker at a given node determined by the diagonal terms of the density matrix.
@@ 3 Network Topologies
In a complete graph, every node is connected to all other nodes, resulting in a homogeneous network, where each node has degree of $N-1$. The cycle graph is also a homogeneous network, where each node is connected only to its two nearest neighbors, yielding a uniform degree of $2$. In contrast, the star graph exhibits a highly heterogeneous structure in which the central hub node has degree $N-1$, while all peripheral (edge) nodes are connected only to the center node and hence have degree $1$.
We consider three types of complex network topologies for our simulations: Barabási–Albert scale-free networks [38], Erdős–Rényi random networks [39], and Watts–Strogatz small-world networks [40], as illustrated in Fig. 1. These network models differ significantly in their structural characteristics due to distinct connectivity patterns.
Erdős–Rényi networks [39] are homogeneous, with a binomial degree distribution tightly centered around the average degree, making the nodes statistically similar in connectivity. In contrast, Barabási–Albert networks [38] are scale-free, characterized by a power-law degree distribution. This results in a few highly connected hub nodes, introducing significant degree heterogeneity into the network.
The Watts–Strogatz model [40] generates small-world networks with high clustering and short average path lengths. Starting from a regular ring lattice where each node is connected to its four nearest neighbors, a small fraction of the edges is randomly rewired with probability $p=0.1$, introducing long-range shortcuts while maintaining local structure [41]. This process disrupts the periodicity of the lattice, creating a network that is partially random but still relatively homogeneous in degree distribution.
@@
4 Stability of quantum networks
@@
4.1 Occupation Probability
We analyze the node occupation probability and its long-time behavior to understand the influence of decoherence on CTQWs. The occupation probabilities corresponds to the diagonal elements of the density matrix. In the presence of decoherence, the occupation probability evolves toward a steady state over time. The rate at which the system approaches its steady state serves as an informative measure of the CTQW’s stability in the presence of decoherence. Slower convergence reflects greater stability of the CTQW under specific decoherence. Moreover, when the occupation probabilities approach uniformity across nodes, the system tends toward classical-like behavior, signaling a transition from quantum to classical [25].
4.2 $\ell_{1}$-norm of coherence
$\ell_{1}$-norm of coherence is the sum of the absolute values of all off-diagonal elements of the density matrix [42].
@@ 4.2 $\ell_{1}$-norm of coherence
4.3 Fidelity
Fidelity [43, 44, 45, 46] quantifies the degree of overlap between the initial quantum state and the state evolved under decoherence. It measures how closely the evolved state overlaps to the initial state. Given the density matrix $\rho(t)$ at time $t$ and the initial density matrix $\rho_{0}$, the fidelity $F$ is defined:
@@ 4.3 Fidelity
4.4 Von Neumann Entropy
Von Neumann entropy measures the degree of mixedness of the quantum state as a result of decoherence [47]. It is defined as:
@@ 4.4 Von Neumann Entropy
4.5 Quantum-Classical Distance
The quantum-classical distance $D_{QC}(t)$ measures the deviation between the quantum evolution and the classical evolution of the system [37]. It is defined as:
@@ 4.5 Quantum-Classical Distance
The quantum-classical distance is a measure that tells the divergence between quantum and classical evolution. If it is 00, then the quantum evolution is fully classicalized, which is not desired, and if $1$, where the system evolves entirely different from classical evolution, indicates that decoherence has not classicalized the system and quantum properties are preserved. Quantum classical distance for complete cycle and star topologies is already studied in [21].
Using these metrics on different topologies, we analyze which topology is more preferred under different decoherences.
5 Results
5.1 CTQW on Complete, Cycle and Star Networks

### v2

@@ Abstract
We investigate the stability of continuous-time quantum walks (CTQWs) in a range of network topologies under different decoherence mechanisms, defining stability as the system’s ability to preserve quantum properties over time. The networks studied range from homogeneous to heterogeneous structures, including cycle, complete, Erdős–Rényi, small-world, scale-free, and star topologies. The decoherence models considered are energy-based intrinsic decoherence, node-based Haken–Strobl noise, and edge-based quantum stochastic walks (QSWs). To assess quantum stability, we employ several metrics: node occupation probabilities, the $\ell_{1}$-norm of coherence, fidelity with the initial state, quantum-classical distance, and von Neumann entropy. Our results reveal that the interplay of both network topology and decoherence model influences coherence preservation. Intrinsic decoherence results in the slowest decay of coherence, followed by Haken–Strobl noise, while QSW causes the most rapid loss of coherence. The stability ranking among network topologies varies depending on the decoherence model and quantifier used. For example, we show that for Haken-Strobl noise, topologies like complete, star, and scale-free with high degree nodes are most stable. Conversely, under QSW decoherence, these same networks with initialization on high degree node becomes uniquely fragile, exhibiting rapid coherence loss. In general, networks such as star and scale-free networks, exhibit the highest stability in all cases except for QSW. However, these same networks, due to their high degree of localization, also show lower values of coherence even in the noiseless case, highlighting a fundamental trade-off between localization and coherence. Furthermore, in heterogeneous networks, the centrality (degree or closeness) of the initialized node has a pronounced impact on stability, underscoring the critical role of local topological features in quantum dynamics.
@@
Keywords: Continuous-time quantum walks, complex networks, decoherence, coherence, stability.
I Introduction
@@
Quantum walks (QWs) are quantum analogs of classical random walks, have emerged as a powerful framework for exploring a wide range of quantum phenomena, including quantum computation, coherent transport, and the dynamical behavior of complex quantum systems [aharonov1993quantum, ambainis2003quantum, konno2008quantum]. Unlike their classical counterparts, QWs exploit quantum superposition and interference, enabling the walker to explore graphs more efficiently and, in some cases, achieve quantum speedups in algorithmic tasks [childs2004spatial].
@@
QWs can be broadly categorized into two types: discrete-time quantum walks (DTQWs) [aharonov1993quantum] and continuous-time quantum walks (CTQWs) [farhi1998quantum]. DTQWs evolve via alternating applications of a coin operator and a conditional shift, allowing fine-grained control over the walker’s trajectory [jayakody2023revisiting]. This approach has proven useful in quantum search algorithms [childs2009universal] and in simulating topologically nontrivial phases [boada2015quantum]. In contrast, CTQWs evolve under a time-independent Hamiltonian defined by the graph structure, without requiring an internal coin degree of freedom. This formulation is particularly well-suited for modeling coherent transport in structured networks and has applications in quantum search [childs2004spatial, chakraborty2016spatial, apers2022quadratic], quantum state transfer [coutinho2025peak], entanglement routing [sauglam2023entanglement], and energy transport in photosynthetic complexes [mohseni2008environment].
@@
However, real-world implementations of QWs are inevitably affected by decoherence arising from environmental interactions and system imperfections. These effects disrupt unitary evolution, suppress quantum coherence, and reduce transport efficiency and computational fidelity [ren2009decoherence, benedetti2016non]. Interestingly, in some cases decoherence can also assist transport [caruso2009highly, plenio2008dephasing]. Understanding how decoherence influences QWs is therefore critical for the development of robust quantum technologies.
@@
Decoherence in open quantum systems is typically continuous rather than discrete [ren2009decoherence], our study focuses on CTQWs under a variety of continuous decoherence mechanisms. Common modeling approaches include intrinsic decoherence, modeled as dephasing in the energy eigenbasis via the Milburn equation [milburn1991intrinsic]; position-basis decoherence, described by the Haken–Strobl master equation [haken1973exactly]; and quantum stochastic walks (QSW), which interpolate between quantum and classical regimes [caruso2014universally]. These frameworks thus enable us to explore different facets of open quantum dynamics and their consequences for quantum transport.
@@
The evolution of CTQWs is influenced by both the form of decoherence and the structure of the underlying graph [bressanini2022decoherence, strauch2009reexamination]. Earlier studies have explored CTQW dynamics on one-dimensional lattices to examine quantum transport [siloi2017noisy] and quantum correlations [benedetti2012quantum], while more recent efforts have extended this work to symmetric and complex networks, assessing how structure interacts with decoherence [bressanini2022decoherence, mulken2011continuous, tsomokos2011quantum].
@@
In this work, we broaden the scope by considering a diverse set of network topologies, and evaluating their behavior under multiple decoherence models. Our primary objective is to understand how the interplay between network structure and environmental noise governs the behavior and stability of CTQWs.
@@
Motivated by analogous findings in classical network theory, where the stability of dynamical states is often determined by the interplay between topology and dynamics [boccaletti2006complex, mulken2011continuous, meena2023emergent], so in this study we seek to uncover similar principles within the context of quantum systems. Specifically, we investigate how the CTQW dynamics evolves under various decoherence mechanisms across different network architectures. Here, we define quantum stability as the ability of a system to retain its quantum features, such as coherence, over time despite the presence of decoherence.
@@
It is well established that node’s centralities, such as degree and closeness, play a critical role in determining the resilience of classical networks [rungta2018identifying, meena2017threshold]. Recent studies indicate that these structural attributes also significantly influence quantum dynamics [böttcher2021classical]. Identifying which topological features help mitigate or exacerbate decoherence is essential for designing stable and efficient quantum systems. Such insights are crucial for advancing robust quantum communication protocols, quantum devices, and fault-tolerant quantum architectures [bose2007quantum]. Motivated by this, our study also investigates how node centralities affect quantum stability, with a particular focus on heterogeneous networks, where centrality disparities are pronounced. To quantify the stability of CTQWs under different decoherence models, we employ several metrics used in the analysis of open quantum systems, including quantum-classical distance, von Neumann entropy, $\ell_{1}$-norm of coherence, and fidelity. These measures provide a multifaceted view of how far the system deviates from ideal quantum behavior over time. For example, the quantum-classical distance has been used to study how CTQWs diverge from quantum behavior under decoherence in simple network topologies such as complete, cycle, and star graphs [bressanini2022decoherence].
@@
We begin Sec. II by detailing the dynamics of CTQWs in the presence of various decoherence mechanisms. In Sec. III, we provide an overview of the network structures considered in our study. Sec. IV introduces the quantitative measures employed to assess the stability of CTQWs across different graph topologies. The primary findings of our investigation are presented in Sec. V, where we analyze how decoherence impacts quantum coherence on both simple and complex networks. Finally, Sec. VI concludes the paper with a summary of our key results and their broader implications.
II Dynamical framework of CTQWs on networks
@@
The Laplacian matrix is defined as L=DAL=D-A, where $D$ is the diagonal degree matrix with the degree (i.e., number of connections) of each node as its diagonal elements, and $A$ is the adjacency matrix, with elements $A_{ij}=1$ if node $i$ and $j$ are connected and $A_{ij}=0$ otherwise.
We investigate the impact of three distinct types of decoherence on CTQW: (i) intrinsic decoherence, modeled by the Milburn equation [milburn1991intrinsic], (ii) position-basis decoherence, described by the Haken–Strobl master equation [haken1973exactly], and (iii) quantum stochastic walks, as formulated in [caruso2014universally]. Each of these models introduces a unique form of non-unitary evolution, enabling us to explore different facets of quantum system dynamics under decoherence on networks.
In the absence of decoherence, that is, ideal noiseless conditions, the quantum walker evolves coherently, governed by unitary dynamics. This evolution is described by the Liouville von Neumann equation [breuer2002theory]:
@@
To model intrinsic decoherence, we take Milburn formalism [milburn1991intrinsic], which introduces non-unitary dynamics by modifying the noiseless evolution Eq. (1) as follows:
@@
Here, $\gamma$ represents the decoherence rate. This model captures intrinsic decoherence, which reflects the system’s internal loss of coherence, independent of external environmental interactions. The double-commuter term in Eq. (2) models decoherence in the energy basis.
The next model we consider is based on the Haken-Strobl master equation, which is commonly employed to describe dephasing noise on the position basis. It takes the following form [haken1973exactly]:
@@
$P_{k}=|k\rangle\langle k|$ are projectors on the position basis. This equation models decoherence by coupling the system to a bath that causes loss of coherence in the position(or node) basis. As the projectors $P_{k}$ act on individual nodes, this is a node based decoherence model.
The last decoherence model is Quantum Stochastic Walk (QSW) [caruso2014universally] that interpolates between pure quantum and classical behavior. It is described by the equation as follows [QSW]:
@@
$P_{kj}=L_{kj}|k\rangle\langle j|$, this jump operators are dependent on connections(edges) between nodes making it a edge based decoherence model. The parameter $p$ controls the interpolation between unitary evolution and stochastic noise.
III Network Topologies
We consider a quantum walker evolving under CTQW dynamics, where the walker is initially localized at a single node in the network. This approach is applied consistently across all the networks that we consider in this study. The rationale behind this choice is to minimize the fidelity between the quantum-evolved state and the classically evolved state, as discussed in prior studies on CTQW [gualtieri2020quantum]. As the system evolves, the walker’s state becomes a quantum superposition over multiple nodes, with the probability of finding the walker at a given node determined by the diagonal terms of the density matrix.
@@ III Network Topologies
In a complete graph, every node is connected to all other nodes, resulting in a homogeneous network, where each node has degree of N1N-1 . The cycle graph is also a homogeneous network, where each node is connected only to its two nearest neighbors, yielding a uniform degree of $2$. In contrast, the star graph exhibits a highly heterogeneous structure in which the central hub node has degree N1N-1, while all peripheral (edge) nodes are connected only to the center node and hence have degree $1$.
We consider three types of complex network topologies for our simulations: Barabási–Albert scale-free networks [barabasi1999emergence], Erdős–Rényi random networks [erdos1960evolution], and Watts–Strogatz small-world networks [watts1998collective], as illustrated in Fig. 1. These network models differ significantly in their structural characteristics due to distinct connectivity patterns.
Erdős–Rényi networks [erdos1960evolution] are homogeneous, with a binomial degree distribution tightly centered around the average degree, making the nodes statistically similar in connectivity. In contrast, Barabási–Albert networks [barabasi1999emergence] are scale-free, characterized by a power-law degree distribution. This results in a few highly connected hub nodes, introducing significant degree heterogeneity into the network.
The Watts–Strogatz model [watts1998collective] generates small-world networks with high clustering and short average path lengths. Starting from a regular ring lattice where each node is connected to its four nearest neighbors, a small fraction of the edges is randomly rewired with probability $p=0.1$, introducing long-range shortcuts while maintaining local structure [mulken2007quantum]. This process disrupts the periodicity of the lattice, creating a network that is partially random but still relatively homogeneous in degree distribution.
@@
IV Stability of CTQW on networks
@@
IV.1 Occupation Probability
We analyze the node occupation probability and its long-time behavior to understand the influence of decoherence on CTQWs. The occupation probabilities corresponds to the diagonal elements of the density matrix. In the presence of decoherence, the occupation probability evolves toward a steady state over time. The rate at which the system approaches its steady state serves as an informative measure of the CTQW’s stability in the presence of decoherence. Slower convergence reflects greater stability of the CTQW under specific decoherence. Moreover, when the occupation probabilities approach uniformity across nodes, the system tends toward classical-like behavior, signaling a transition from quantum to classical [mulken2011continuous].
IV.2 $\ell_{1}$-norm of coherence
$\ell_{1}$-norm of coherence is the sum of the absolute values of all off-diagonal elements of the density matrix [baumgratz2014quantifying].
@@ IV.2 $\ell_{1}$-norm of coherence
IV.3 Fidelity
Fidelity [jozsa1994fidelity, raginsky2001fidelity, gilchrist2005distance, chen2018simulating] quantifies the degree of overlap between the initial quantum state and the state evolved under decoherence. It measures how closely the evolved state overlaps to the initial state. Given the density matrix $\rho(t)$ at time $t$ and the initial density matrix $\rho(t)$, the fidelity $F$ is defined:
@@ IV.3 Fidelity
IV.4 Von Neumann Entropy
Von Neumann entropy measures the degree of mixedness of the quantum state as a result of decoherence [rieffel2011quantum]. It is defined as:
@@ IV.4 Von Neumann Entropy
IV.5 Quantum-Classical Distance
The quantum-classical distance $D_{QC}(t)$ measures the deviation between the quantum evolution and the classical evolution of the system [gualtieri2020quantum]. It is defined as:
@@ IV.5 Quantum-Classical Distance
The quantum-classical distance is a measure that tells the divergence between quantum and classical evolution. If it is 0, then the quantum evolution is fully classicalized, which is not desired, and if $1$, where the system evolves entirely different from classical evolution, indicates that decoherence has not classicalized the system and quantum properties are preserved. Quantum classical distance for complete cycle and star topologies is already studied in [bressanini2022decoherence].
Using these metrics on different topologies, we analyze which topology is more preferred under different decoherence.
V Results
V.1 CTQW on Complete, Cycle and Star Networks
Figure 2: Effect of network size and decoherence rate on the $\ell_{1}$-norm of coherence for a star graph with central-hub initialization. Panels (a)–(c) show the coherence decay for different network sizes NN under (a) Haken–Strobl noise, (b) QSW dynamics, and (c) intrinsic decoherence, each evaluated at fixed decoherence rates ($\gamma=0.1$ or $p=0.1$). (d)–(f) The bottom row illustrates the coherence decay for increasing decoherence rates ($\gamma$ or $p$) for the three corresponding decoherence models on a network of fixed size $N=10$.
No noiseHaken-StroblIntrinsic DecoherenceQuantum StochasticRefer to caption